OpenStudy (anonymous):

how do you find the partial fraction of: 1/[(s^2+1)(s^2+s+5/4)] ?

OpenStudy (dumbcow):

\[\frac{As+B}{s^{2} +1}+\frac{Cs+D}{s^{2} +s+5/4}\] \[A+C = 0\] \[A+B+D =0\] \[(5/4)A +B+C = 0\] \[(5/4)B +D = 1\]

OpenStudy (anonymous):

i got it. thanks!